# Generic free fermions with nearest neighbour interactions

### v1

@@ Acknowledgements.
We thank Maurizio Fagotti for useful discussions. We acknowledge financial support from the Royal Society through the University Research Fellowship No. 201101 and warmly acknowledge the hospitality of the Simons Center for Geometry and Physics during the program “Fluctuations, Entanglements, and Chaos: Exact Results” where part of the project has been carried out.
Appendix A Higher order harmonics
In this appendix we discuss how the inclusion of higher harmonics helps achieving rational independence also for non-prime sizes.
First we note that our counter example from Sec. II does not apply if one adds higher harmonics. To see this consider again $L=np$ with $p$ prime and add the term
$P(1)=0,$ $P(\xi)=0,$ $P(\xi^{p})=0.$ (42)
Second, we show that when $n$ is also prime we have rational independence if we add harmonics of order $p$ and $n$. This can be seen as follows. Assuming again incommensurability, we obtain that (11) can be decomposed into the following four equations
$P(1)=0,$ $P(\xi)=0,$ $P(\xi^{p})=0,$ $P(\xi^{n})=0.$ (43)
$P(z)=R_{1}(z)\Phi_{np}(z)=R_{2}(z)\Phi_{n}(z)=R_{3}(z)\Phi_{p}(z).$ (44)
$P(z)=R_{4}(z)\Phi_{n}(z)\Phi_{p}(z)\Phi_{np}(z).$ (45)
for some $R_{4}(z)$.
$\prod_{d|n}\Phi_{d}(z)=z^{n}-1.$ (46)
This tells us that $\Phi_{n}(z)\Phi_{p}(z)\Phi_{np}(z)$ is by definition
$\Phi_{n}(z)\Phi_{p}(z)\Phi_{np}(z)=1+z+\dots+z^{pn-1},$ (47)

### v2

@@ Acknowledgements.
We thank Maurizio Fagotti and Hosho Katsura for useful discussions. We thank Henrik Wilming and Lorenzo Campos Venuti for bringing Refs. Shiraishi and Tasaki (2024) and Tasaki (2010) to our attention. We also thank Lorenzo Campos Venuti for pointing out a mistake in Eq. (49) of the first version of this paper. We acknowledge financial support from the Royal Society through the University Research Fellowship No. 201101 and warmly acknowledge the hospitality of the Simons Center for Geometry and Physics during the program “Fluctuations, Entanglements, and Chaos: Exact Results” where part of the project has been carried out.